Mary's risk aversion is 2.1. What percent of her savings should she invest in a portfolio with E(r)=18% and standard deviation of 20%, if the risk free rate to invest in is 3.7% and the rate at which money can be borrowed is 5%? Provide your answer in percent, rounded to two decimals, omitting the % sign.
Answer #1

Std dev = 0.2

Variance = 0.2 * 0.2 = 0.04

U = E(r) – 0.5 x A x σ2

= 0.18 - 0.5 * 2.1 * 0.04

= 0.138 or 13.8%

Since this utility rate of 0.138 is much higher than the risk free rate of 3.7%, she should definitely invest in the portfolio.